# Fairness-aware Contextual Dynamic Pricing with Strategic Buyers

### v1

@@ 5.3 Sensitivity Tests
6 Real Application
In this section, we evaluate the efficacy of our policy using the public Home Mortgage Disclosure Act (HMDA) dataset that includes customer characteristics and loan features. In our study, we employ the 2022 HMDA dataset [^9] https://ffiec.cfpb.gov/data-publication/dynamic-national-loan-level-dataset/2022. HMDA mandates that numerous financial institutions maintain, report, and publicly disclose mortgage-related information. This dataset has been a focal point of prior research (Bocian et al., 2008; Zhang, 2018; Bartlett et al., 2022; Popick, 2022; Butler et al., 2023), revealing that borrowers from minority groups often face higher interest rates even after accounting for contextual factors.
6.1 Data Description and Preprocessing
We only retain data for loans that have been approved and disbursed to borrowers. We designate race as the group status, loan amount as the demand, interest rate as the price, and contextual information comprises income, age, property value securing the loan, debt-to-income ratio, combined loan-to-value ratio, and loan term. Race is categorized into two groups: White and non-White, encompassing Black or African American, American Indian or Alaska Native, Native Hawaiian or Other Pacific Islander, Asian and other minority races. We consider borrowers aged between 25 and 74. Initially provided in discrete intervals, we transform age into continuous data by averaging within each interval. The debt-to-income ratio is limited to the range of 20% to 60%. To ensure data integrity, we eliminate outliers, excluding the upper 5% and lower 5% of values, for loan amount, interest rate, income, property value, combined loan-to-value ratio, and loan term. Following data preprocessing, our dataset comprises 3,871,912 records. Among them, the White group constitutes 81.86% of the dataset.
6.2 Data Analysis
Before applying our pricing policy on the dataset, we employ the propensity score matching method (Ho et al., 2007, 2011) to examine the interest rate difference between the White group and the non-White group after controlling the contextual information. Let $\bar{p}_{0}$ and $\bar{p}_{1}$ be the mean prices of the non-White group and White group, respective. We test the hypothesis
$H_{0}:\bar{p}_{1}=\bar{p}_{0},\ H_{1}:\bar{p}_{1}<\bar{p}_{0}.$ (9)
The $p$-value of the hypothesis problem (9) is $2.2\times 10^{-16}$, indicating that the non-White group pays a higher interest rate compared to their White counterparts. This finding aligns with previous research (Bocian et al., 2008; Zhang, 2018; Bartlett et al., 2022; Popick, 2022; Butler et al., 2023).
Figure 7: Regret plots for the two policies on the real data. The remaining caption is the same as Figure 2.
Now, we analyze the dataset using our online pricing policy. In practice, obtaining real-time feedback from buyers regarding any dynamic pricing strategy is challenging until the pricing policy has been executed in the data collection system. Consequently, we adhere to the calibration approach outlined in previous studies (Fan et al., 2024; Wang et al., 2024; Liu et al., 2024) by initially estimating a linear demand model. This model serves as a ground truth to evaluate dynamic pricing policies. We set the manipulation cost at $C_{0}=0.11$ and the fairness threshold to $\delta=0.1$. We assume $\epsilon_{t}\sim N(0,1)$, and fix $B=1$ and $\tau=5$. Buyers learn the price difference using a neural network. Our policy is applied with the time horizon $T=50000$ , repeated 20 times to record average cumulative regrets. We compare our policy against the benchmark pricing policy without considering buyers’ fairness learning. Figure 7 shows that the cumulative regret of the policy without buyers’ learning is much larger and grows significantly faster compared to our pricing policy, aligning with our earlier findings in simulated data. Our proposed pricing policy achieves 35.06% reduction in regret compared to the benchmark policy at the end of the time horizon.
7 Summary and Future Directions
Looking ahead, there are several promising avenues for future exploration in this area. We can extend our investigation to incorporate additional complexities such as strategic pricing problems with censored demand (Tang et al., 2025), unobserved confounding (Yu et al., 2022; Miao et al., 2023; Qi et al., 2024; Shi et al., 2024), offline learning (Duan et al., 2021; Duan and Wainwright, 2024) and other fairness constraints (Fang et al., 2023). Exploring these avenues will aid in developing more robust and equitable dynamic pricing strategies for online retail environments.
Supplementary Materials
“Fairness-aware Contextual Dynamic Pricing with Strategic Buyers”
[…]

### v2

@@ Acknowledgment
Appendix S.2 Estimation Errors of Neural Network and Decision Tree
Figure 8 illustrates the estimation errors of $\delta$ under the neural network and the decision tree, showing that the decision tree yields higher error. These higher errors reduce its effectiveness in discouraging strategic manipulation, which in turn leads to greater regret as shown in Figure 2 of Section 5.1.
Refer to caption Refer to caption
Figure 8: Estimation errors of neural network and decision tree.
Appendix S.3 Additional Details of Real Data Analysis
We first describe the potential misreporting mechanism in the real application. During the load application process, race can be self-identified at the time of application; however, when the information is missing, lenders are permitted to record race based on visual observation or surname. Our notion of “misreporting” does not assume that the borrower intentionally provides false information at the application stage. Instead, we focus on a potential mismatch that can arise through the appraisal process. In the real application, the borrower already owns the property being appraised and applies for a loan for another purpose. The lender orders an appraisal, and the appraiser conducts an in-person inspection of the property. During this visit, the appraiser typically meets an occupant who provides access to the home and may answer questions about the property. The appraiser does not verify legal ownership or the identity of the borrower. Their task is to assess the physical characteristics and market value of the property. Because the appraiser may rely on visual cues about the person present at the inspection, a mismatch can occur if that person is not the borrower. For example, if the borrower asks a “friend" of a different race to be present to grant access. In that case, the appraiser’s perception of the “homeowner” can differ from the borrower’s true race. Our model uses this type of appraisal-visit context as a channel through which the perceived race can diverge from the borrower’s true race, thereby generating the misreporting mechanism we analyze.
We next use the R package MatchIt (Ho et al., 2011) to perform propensity score matching. Propensity scores are estimated via the logistic regression, and we apply 4-to-1 nearest neighbor matching—that is, for each borrower in the non-White group, we match the four closest White borrowers based on propensity scores.
Table 2: Summary of Covariate Balance Before and After Matching
Before Matching After Matching
Covariate Mean Diff Var Ratio eCDF Mean Diff Var Ratio eCDF
Income 0.0364 0.9932 0.007 0.0305 0.9988 0.0059
Loan term 0.0444 0.8685 0.0098 0.0379 0.8867 0.0084
Debt/income 0.2151 0.9925 0.0705 0.183 0.9809 0.0597
Loan/value 0.2303 0.8800 0.0563 0.2006 0.8874 0.0490
Property value 0.1328 1.0244 0.0264 0.1083 1.0064 0.0216
Discount points 0.1895 1.4118 0.0572 0.1478 1.2029 0.0473
Lender credits 0.1198 1.0930 0.0380 0.0844 0.7006 0.0335
Applicant age 0.0178 0.9082 0.0133 0.0142 0.9085 0.0128
Loan purpose 1 0.1634 - 0.0732 0.1440 - 0.0645
Loan purpose 2 -0.0274 - 0.0012 -0.0248 - 0.0011
Loan purpose 3 -0.1064 - 0.0239 -0.0944 - 0.0212
Loan purpose 4 -0.1155 - 0.0480 -0.1014 - 0.0422
State code 1 -0.0093 - 0.0004 -0.0066 - 0.0003
State code 2 0.0142 - 0.0020 0.0201 - 0.0028
State code 3 -0.0514 - 0.0033 -0.0455 - 0.0029
State code 4 -0.0964 - 0.0139 -0.0831 - 0.0120
State code 5 0.0901 - 0.0253 0.0827 - 0.0233
State code 6 -0.1402 - 0.0162 -0.1313 - 0.0152
State code 7 -0.0081 - 0.0007 -0.002 - 0.0002
State code 8 0.0363 - 0.0018 0.0254 - 0.0013
State code 9 0.0250 - 0.0019 0.0236 - 0.0018
State code 10 -0.0659 - 0.0191 -0.0424 - 0.0123
State code 11 0.2028 - 0.0597 0.1521 - 0.0448
State code 12 0.0128 - 0.0002 0.0098 - 0.0001
State code 13 0.0699 - 0.0056 0.0456 - 0.0037
State code 14 -0.0969 - 0.0038 -0.0918 - 0.0036
State code 15 -0.1714 - 0.0059 -0.1587 - 0.0055
State code 16 0.0252 - 0.0041 0.0305 - 0.0049
State code 17 -0.0622 - 0.0076 -0.0507 - 0.0062
State code 18 -0.0583 - 0.0031 -0.0542 - 0.0028
State code 19 -0.0983 - 0.0071 -0.0934 - 0.0067
State code 20 0.0407 - 0.0045 0.037 - 0.0041
State code 21 -0.0114 - 0.0012 -0.0044 - 0.0005
State code 22 0.1446 - 0.0317 0.1073 - 0.0235
State code 23 -0.1479 - 0.0035 -0.1358 - 0.0032
State code 24 -0.1210 - 0.0140 -0.111 - 0.0129
State code 25 -0.0917 - 0.0107 -0.081 - 0.0095
State code 26 -0.1052 - 0.0087 -0.099 - 0.0082
State code 27 0.0256 - 0.0021 0.0248 - 0.0020
State code 28 -0.2259 - 0.0049 -0.205 - 0.0045
State code 29 0.0721 - 0.0160 0.0675 - 0.0150
State code 30 -0.0624 - 0.0013 -0.0598 - 0.0012
State code 31 -0.0716 - 0.0028 -0.0684 - 0.0027
State code 32 -0.1522 - 0.0039 -0.1393 - 0.0036
State code 33 0.0221 - 0.0036 0.0258 - 0.0042
State code 34 -0.1001 - 0.0050 -0.0957 - 0.0048
State code 35 0.0412 - 0.0058 0.0382 - 0.0054
State code 36 0.0631 - 0.0123 0.0583 - 0.0114
State code 37 -0.1076 - 0.0159 -0.0933 - 0.0137
State code 38 -0.0311 - 0.0031 -0.0227 - 0.0023
State code 39 -0.1197 - 0.0100 -0.1135 - 0.0095
State code 40 -0.0612 - 0.0091 -0.0470 - 0.0070
State code 41 -0.1108 - 0.0025 -0.1032 - 0.0023
State code 42 -0.0351 - 0.0014 -0.0315 - 0.0013
State code 43 0.0034 - 0.0005 0.0117 - 0.0017
State code 44 -0.1022 - 0.0021 -0.0943 - 0.0020
State code 45 -0.0919 - 0.0129 -0.0795 - 0.0112
State code 46 0.1109 - 0.0392 0.1077 - 0.0381
State code 47 -0.1815 - 0.0120 -0.1721 - 0.0113
State code 48 0.0643 - 0.0126 0.0586 - 0.0115
State code 49 -0.0026 - 0 -0.0023 - 0
State code 50 -0.0905 - 0.0009 -0.0828 - 0.0008
State code 51 -0.0427 - 0.0062 -0.0313 - 0.0046
State code 52 -0.1491 - 0.0108 -0.1432 - 0.0104
State code 53 -0.0751 - 0.0026 -0.0712 - 0.0024
State code 54 -0.1427 - 0.0021 -0.1292 - 0.0019
Table 2: Summary of Covariate Balance Before and After Matching (Continued)
Table LABEL:tab2 summarizes the covariate balance before and after matching. Balance is evaluated using three metrics: standardized mean differences (Mean Diff), variance ratios (Var Ratio), and the mean distance between the two empirical quantile functions (eCDF). The variance ratios for the one-hot encoded variables are not available. In general, values of Mean Diff and eCDF close to zero, and Var Ratio close to one, indicate good balance. As seen in the “Before Matching” columns, covariates exhibit notable imbalance. After matching, balance improves substantially for most covariates.
The model fits the data well for both groups. The $R^{2}$ value for the non-White group is 0.966, and for the White group, it is 0.960. We provide the estimated regression coefficients for both groups in Tables LABEL:tab:nonwhite-regression and LABEL:tab:white-regression. In the tables, “Debt/income" represents the debt-to-income ratio, and “Loan/value" represents the combined loan-to-value ratio. All covariates are statistically significant. The estimated coefficients are also intuitively reasonable. For example, the coefficient on interest rate is negative, implying that higher rates reduce loan demand. The coefficient on applicant age is negative, suggesting that younger individuals are more likely to apply for larger loan amounts.
Table 3: Regression Estimates for Non-White Group
Covariate Coef Std Err $t$ $P>|t|$ 95% confidence interval
Constant 24.3732 2.096 11.628 0 (20.265, 28.481)
Interest rate -0.746 0.086 -8.636 0 (-0.915, -0.577)
Income 6.0192 0.109 55.012 0 (5.805, 6.234)
Loan term -0.5446 0.09 -6.029 0 (-0.722, -0.368)
Debt/income 1.4636 0.091 16.055 0 (1.285, 1.642)
Loan/value 62.93 0.116 543.503 0 (62.703, 63.157)
Property value 132.6207 0.122 1088.608 0 (132.382, 132.86)
Discount points 1.834 0.078 23.543 0 (1.681, 1.987)
Lender credits 0.6738 0.08 8.435 0 (0.517, 0.830)
Applicant age -1.0056 0.09 -11.22 0 (-1.181, -0.830)
Loan purpose 2 3.8792 1.829 2.121 0.034 (0.294, 7.464)
Loan purpose 3 2.3553 0.398 5.92 0 (1.576, 3.135)
Loan purpose 4 7.0774 0.245 28.882 0 (6.597, 7.558)
State code 2 -9.0071 2.128 -4.232 0 (-13.179, -4.835)
State code 3 -10.368 2.398 -4.324 0 (-15.068, -5.668)
State code 4 2.2939 2.123 1.08 0.28 (-1.868, 6.456)
State code 5 -2.1351 2.073 -1.03 0.303 (-6.198, 1.927)
State code 6 5.3019 2.165 2.448 0.014 (1.058, 9.546)
State code 7 -3.1402 2.244 -1.4 0.162 (-7.538, 1.257)
State code 8 -10.4012 2.602 -3.997 0 (-15.502, -5.300)
State code 9 -2.5258 2.295 -1.1 0.271 (-7.025, 1.973)
State code 10 -0.7204 2.068 -0.348 0.728 (-4.773, 3.332)
State code 11 -0.9517 2.067 -0.46 0.645 (-5.003, 3.100)
State code 12 5.5142 6.021 0.916 0.36 (-6.288, 17.316)
State code 13 -11.5147 2.282 -5.045 0 (-15.988, -7.041)
State code 14 -8.3882 2.883 -2.91 0.004 (-14.038, -2.738)
State code 15 -3.2503 3.09 -1.052 0.293 (-9.307, 2.807)
State code 16 -7.8519 2.108 -3.725 0 (-11.984, -3.720)
State code 17 -3.6947 2.152 -1.717 0.086 (-7.913, 0.523)
State code 18 -6.73 2.555 -2.634 0.008 (-11.737, -1.723)
State code 19 -5.7833 2.33 -2.482 0.013 (-10.350, -1.216)
State code 20 -13.7401 2.175 -6.317 0 (-18.003, -9.477)
State code 21 -0.5792 2.182 -0.265 0.791 (-4.857, 3.698)
State code 22 0.376 2.082 0.181 0.857 (-3.704, 4.456)
State code 23 4.1571 3.951 1.052 0.293 (-3.587, 11.901)
State code 24 -7.9109 2.162 -3.659 0 (-12.149, -3.673)
State code 25 0.1378 2.16 0.064 0.949 (-4.096, 4.371)
State code 26 -7.0645 2.264 -3.12 0.002 (-11.502, -2.627)
State code 27 -9.3169 2.269 -4.107 0 (-13.764, -4.870)
State code 28 -2.5528 4.202 -0.607 0.544 (-10.790, 5.684)
State code 29 -0.2927 2.081 -0.141 0.888 (-4.371, 3.786)
State code 30 -3.8272 4.416 -0.867 0.386 (-12.482, 4.827)
State code 31 -2.9708 2.883 -1.031 0.303 (-8.621, 2.679)
State code 32 3.8654 3.718 1.04 0.298 (-3.421, 11.152)
State code 33 -0.046 2.109 -0.022 0.983 (-4.179, 4.087)
State code 34 -2.2117 2.605 -0.849 0.396 (-7.317, 2.893)
State code 35 1.8927 2.127 0.89 0.374 (-2.276, 6.061)
State code 36 -5.4352 2.092 -2.598 0.009 (-9.536, -1.334)
State code 37 -6.2738 2.12 -2.959 0.003 (-10.429, -2.118)
State code 38 -10.3223 2.2 -4.691 0 (-14.635, -6.010)
State code 39 2.2401 2.263 0.99 0.322 (-2.195, 6.675)
State code 40 -5.3325 2.118 -2.517 0.012 (-9.485, -1.181)
State code 41 -19.0629 4.142 -4.603 0 (-27.181, -10.945)
State code 42 2.7974 2.845 0.983 0.325 (-2.779, 8.373)
State code 43 -4.1728 2.118 -1.97 0.049 (-8.325, -0.021)
State code 44 -3.5824 4.339 -0.826 0.409 (-12.087, 4.922)
State code 45 2.4697 2.127 1.161 0.246 (-1.698, 6.638)
State code 46 -0.2248 2.062 -0.109 0.913 (-4.267, 3.817)
State code 47 5.6354 2.384 2.364 0.018 (0.963, 10.308)
State code 48 -0.8347 2.089 -0.399 0.69 (-4.930, 3.261)
State code 49 -8.10 1.96 -4.135 0 (-11.900, -4.260)
State code 50 -3.0631 8.085 -0.379 0.705 (-18.910, 12.784)
State code 51 4.6413 2.124 2.185 0.029 (0.478, 8.804)
State code 52 -10.5734 2.326 -4.545 0 (-15.133, -6.014)
State code 53 -5.8529 3.118 -1.877 0.061 (-11.964, 0.258)
State code 54 -6.2122 5.784 -1.074 0.283 (-17.549, 5.125)
Note: Coef, Std Err and 95% confidence interval for “State code 49” are scaled by $10^{-15}$.
Table 3: Regression Estimates for Non-White Group (Continued)
Table 4: Regression Estimates for White Group
Covariate Coef Std Err $t$ $P>|t|$ 95% confidence interval
Constant 24.4108 1.03 23.698 0 (22.392, 26.430)
Interest rate -0.8865 0.047 -18.859 0 (-0.979, -0.794)
Income 5.9661 0.06 99.847 0 (5.849, 6.083)
Loan term -0.9515 0.046 -20.504 0 (-1.043, -0.861)
Debt/income 1.7647 0.05 35.603 0 (1.668, 1.862)
Loan/value 58.7266 0.06 985.327 0 (58.610, 58.843)
Property value 126.8843 0.067 1897.579 0 (126.753, 127.015)
Discount points 2.5833 0.05 51.225 0 (2.484, 2.682)
Lender credits 0.7625 0.045 16.9 0 (0.674, 0.851)
Applicant age -0.8327 0.046 -18.005 0 (-0.923, -0.742)
Loan purpose 2 2.7612 0.784 3.522 0 (1.225, 4.298)
Loan purpose 3 0.5968 0.184 3.236 0.001 (0.235, 0.958)
Loan purpose 4 4.0067 0.121 33.198 0 (3.770, 4.243)
State code 2 -4.2735 1.056 -4.046 0 (-6.344, -2.203)
State code 3 -6.4937 1.124 -5.776 0 (-8.697, -4.290)
State code 4 1.5555 1.031 1.509 0.131 (-0.465, 3.576)
State code 5 -0.2199 1.021 -0.215 0.829 (-2.220, 1.781)
State code 6 4.527 1.036 4.368 0 (2.496, 6.558)
State code 7 -2.4666 1.111 -2.221 0.026 (-4.643, -0.290)
State code 8 2.5675 1.943 1.321 0.186 (-1.241, 6.376)
State code 9 -2.0066 1.214 -1.653 0.098 (-4.385, 0.372)
State code 10 -0.3086 1.013 -0.305 0.761 (-2.294, 1.676)
State code 11 -1.3342 1.03 -1.295 0.195 (-3.353, 0.685)
State code 12 13.1661 10.148 1.297 0.194 (-6.723, 33.055)
State code 13 -2.2878 1.778 -1.287 0.198 (-5.773, 1.197)
State code 14 -6.9136 1.167 -5.924 0 (-9.201, -4.626)
State code 15 -1.2245 1.13 -1.084 0.278 (-3.439, 0.990)
State code 16 -5.4549 1.045 -5.221 0 (-7.503, -3.407)
State code 17 -4.641 1.046 -4.438 0 (-6.690, -2.592)
State code 18 -7.3294 1.156 -6.342 0 (-9.595, -5.064)
State code 19 -5.6874 1.079 -5.273 0 (-7.801, -3.573)
State code 20 -8.3807 1.12 -7.484 0 (-10.576, -6.186)
State code 21 -1.5276 1.076 -1.42 0.156 (-3.636, 0.581)
State code 22 0.7522 1.053 0.715 0.475 (-1.311, 2.816)
State code 23 -2.4104 1.215 -1.984 0.047 (-4.792, -0.029)
State code 24 -5.3938 1.038 -5.195 0 (-7.429, -3.359)
State code 25 -0.1304 1.042 -0.125 0.9 (-2.172, 1.911)
State code 26 -4.4909 1.063 -4.224 0 (-6.574, -2.407)
State code 27 -7.8866 1.189 -6.633 0 (-10.217, -5.556)
State code 28 -2.7618 1.167 -2.367 0.018 (-5.048, -0.475)
State code 29 -1.2667 1.03 -1.23 0.219 (-3.286, 0.753)
State code 30 -7.084 1.459 -4.855 0 (-9.944, -4.224)
State code 31 -3.3698 1.201 -2.806 0.005 (-5.724, -1.016)
State code 32 0.4483 1.193 0.376 0.707 (-1.890, 2.786)
State code 33 -0.0129 1.044 -0.012 0.99 (-2.060, 2.034)
State code 34 -3.5158 1.124 -3.128 0.002 (-5.719, -1.313)
State code 35 2.5048 1.068 2.346 0.019 (0.412, 4.597)
State code 36 -3.6208 1.038 -3.488 0 (-5.656, -1.586)
State code 37 -7.2239 1.029 -7.019 0 (-9.241, -5.207)
State code 38 -7.4766 1.074 -6.963 0 (-9.581, -5.372)
State code 39 1.2463 1.058 1.177 0.239 (-0.828, 3.321)
State code 40 -5.2714 1.034 -5.1 0 (-7.297, -3.246)
State code 41 -17.3413 1.286 -13.483 0 (-19.862, -14.82)
State code 42 0.6704 1.276 0.525 0.599 (-1.831, 3.171)
State code 43 -1.1719 1.046 -1.12 0.263 (-3.222, 0.878)
State code 44 -1.4613 1.322 -1.106 0.269 (-4.052, 1.129)
State code 45 -0.1416 1.032 -0.137 0.891 (-2.165, 1.882)
State code 46 -3.5493 1.013 -3.503 0 (-5.535, -1.564)
State code 47 2.8721 1.062 2.705 0.007 (0.791, 4.953)
State code 48 0.8482 1.038 0.817 0.414 (-1.187, 2.883)
State code 49 1.3431 18.918 0.071 0.943 (-35.735, 38.421)
State code 50 -2.0441 1.689 -1.21 0.226 (-5.355, 1.267)
State code 51 1.3177 1.038 1.269 0.204 (-0.717, 3.352)
State code 52 -4.926 1.061 -4.641 0 (-7.006, -2.846)
State code 53 -9.3132 1.233 -7.554 0 (-11.730, -6.897)
State code 54 -2.0485 1.351 -1.517 0.129 (-4.696, 0.599)
Table 4: Regression Estimates for White Group (Continued)
To assess robustness, we conduct a sensitivity analysis using a milder trimming rule that removes only the upper and lower 1% of selected covariates. The resulting regret curves (Figure 9) are qualitatively similar to those obtained under 5% trimming across all three scenarios ($\delta\in{0.05,0.1,0.2}$), indicating that our main conclusions do not depend on the specific trimming threshold.
Refer to caption Refer to caption Refer to caption
Figure 9: Regret plots for the two policies (1% tails removed). The three subplots show the regrets of three different scenarios, $\delta\in\{0.05,0.1,0.2\}$.
Appendix S.4 Misspecified Demand Models
In this section, we conducted additional simulation experiments to evaluate the performance of our proposed pricing policy under misspecified demand models. The experimental setup follows the same framework as in Section 5, with modifications to the true underlying demand structure. We consider the following two cases of model misspecification:
$\text{Case 1}:y_{j}=\alpha_{j}p+\bm{\beta}_{j0}+\bm{\beta}_{j1}\bm{x}_{1}^{2}+\bm{\beta}_{j2}\bm{x}_{2}+\bm{\beta}_{j3}\bm{x}_{3}+\epsilon.$
and
$\text{Case 2}:y_{j}=\alpha_{j}p+\bm{\beta}_{j0}+\bm{\beta}_{j1}\bm{x}_{1}^{2}+\bm{\beta}_{j2}\bm{x}_{1}\bm{x}_{2}+\bm{\beta}_{j3}\bm{x}_{3}+\epsilon.$
In both settings, the seller incorrectly assumes a linear specification and learns the model using $\widehat{y}_{j}=\widehat{\alpha}_{j}p+\widehat{\bm{\beta}}_{j0}+\widehat{\bm{\beta}}_{j1}\bm{x}_{1}+\widehat{\bm{\beta}}_{j2}\bm{x}_{2}+\widehat{\bm{\beta}}_{j3}\bm{x}_{3}$. Under the misspecification, we conduct simulations to compare the performance of our proposed policy with the benchmark policy. The results are presented in Figure 10 and Figure 11. As expected, the regret under the misspecified setting is higher than in the correctly specified case. The regret in Case 2 is higher than that in Case 1 because the degree of model misspecification is more severe in Case 2. This increased mismatch between the true demand model and the model used by the policy leads to less accurate parameter estimation and higher cumulative regret. However, our policy still significantly outperforms the benchmark in all scenarios.
Refer to caption Refer to caption Refer to caption
Figure 10: Regret plots for the two policies under the misspecified demand model. The three subplots show the regrets of three different scenarios, $c_{\delta}\in\{1,2,3\}$.
Refer to caption Refer to caption Refer to caption
Figure 11: Regret plots for the two policies under the misspecified demand model. The three subplots show the regrets of three different scenarios, $c_{\delta}\in\{1,2,3\}$.
Appendix S.5 Extension to Multi-group Settings
We outline how the proposed framework can be extended to a multi-group setting. Assume there are $N\geq 2$ groups. At time $t$, the demand of a buyer with feature $\bm{x}_{t}$ and group status $G_{t}=j\in\{1,\cdots,N\}$ is $y_{jt}=\alpha_{j}p_{t}+\bm{\beta}_{j}^{\top}\widetilde{\bm{x}}_{t}+\epsilon_{t}$, where $\alpha_{j}$ and $\bm{\beta}_{j}$ are group-specific unknown parameters, $\widetilde{\bm{x}}_{t}=(1,\bm{x}_{t}^{\top})^{\top}$, and $\epsilon_{t}$ is an $i.i.d.$ sub-Gaussian variable with $\mathbb{E}(\epsilon_{t}|\bm{x}_{t},p_{t})=0$. Let $q_{j}$ denote the proportion of buyers from group $j$. At each time $t$, the fairness-aware clairvoyant seller maximizes the weighted revenue by solving the following constrained optimization problem,
$\mathop{\max}_{p_{1},\cdots,p_{N}}$ $\ \sum_{j=1}^{N}q_{j}R_{j}(p_{j},\bm{x}_{t})$ (S1)
$s.t.\ \lvert p_{i}-p_{j}\rvert\leq\delta,\forall i\neq j.$
where $R_{j}(p,\bm{x}_{t})=p(\alpha_{j}p+\bm{\beta}_{j}^{\top}\widetilde{\bm{x}}_{t})$ is the expected revenue. The fairness constraint used here is similar to that proposed in Cohen et al. (2025). We denote $\bm{\theta}_{j}=(\alpha_{j},\bm{\beta}_{j}^{\top})^{\top}$ and $\bm{\theta}=(\bm{\theta}_{1}^{\top},\cdots,\bm{\theta}_{N}^{\top})$. Solving the optimization problem (S1) yields the optimal price for group $j$: $p^{*}_{j}(\bm{\theta},\bm{x}_{t},\delta)$ for $j=1,\cdots,N$. As in the two-group case, the pricing policy consists of an exploration phase and an exploitation phase. During exploration, the seller uses a uniform pricing policy to collect data for estimating the demand parameters $\bm{\theta}$, resulting in estimates $\widehat{\bm{\theta}}$. In the exploitation phase, the seller applies the fairness-aware pricing policy: $p_{j}(\widehat{\bm{\theta}},\bm{x}_{t},\delta_{j})$ for $j=1,\cdots,N$, where $\delta_{j}$ reflects the fairness constraint associated with group $j$. The pricing policy $p_{j}(\widehat{\bm{\theta}},\bm{x}_{t},\delta_{j})$ can be viewed as a plug-in estimator of $p^{*}_{j}(\bm{\theta},\bm{x}_{t},\delta)$, where $\widehat{\bm{\theta}}_{j}$ is obtained from the exploration phase. The fairness level $\delta_{j}$ should typically be chosen close to the global fairness parameter $\delta$, but its exact value must be carefully selected to effectively deter strategic misreporting. Our strategy to select the fairness level in the algorithm can be extended to this multi-group setting.
Appendix S.6 One-time Fixed Cost
If we treat the manipulation cost $C_{0}$ as a one-time fixed amount, the model can be adjusted accordingly. Following our original model, we consider that the group 0 is discriminated. For a buyer from group 0, the expected benefit of manipulation is
$p_{0}(\alpha_{0}p_{0}+\bm{\beta}_{0}^{\top}\widetilde{\bm{x}}_{t})-p_{1}(\alpha_{0}p_{1}+\bm{\beta}_{0}^{\top}\widetilde{\bm{x}}_{t})=\alpha_{0}(p_{0}^{2}-p_{1}^{2})+(p_{0}-p_{1})\bm{\beta}_{0}^{\top}\widetilde{\bm{x}}_{t}.$
To deter misreporting, the seller must ensure this gain does not exceed the fixed manipulation cost $C_{0}$. Accordingly, the fairness-constrained pricing problem becomes
$\mathop{\min}_{p_{0},p_{1}}$ $\ -qR_{0}(p_{0},\bm{x}_{t})-(1-q)R_{1}(p_{1},\bm{x}_{t})$ (S2)
$s.t.\ \alpha_{0}(p_{0}^{2}-p_{1}^{2})+(p_{0}-p_{1})\bm{\beta}_{0}^{\top}\widetilde{\bm{x}}_{t}\leq\delta,$
where $R_{j}(p,\bm{x}_{t})=p(\alpha_{j}p+\bm{\beta}_{j}^{\top}\widetilde{\bm{x}}_{t})$ is the expected revenue and $\delta<C_{0}$. We denote $\bm{\theta}_{j}=(\alpha_{j},\bm{\beta}_{j}^{\top})^{\top}$ and $\bm{\theta}=(\bm{\theta}_{0}^{\top},\bm{\theta}_{1}^{\top})$. Solving the optimization problem (S2) yields the optimal price for group $j$: $p^{*}_{j}(\bm{\theta},\bm{x}_{t},\delta)$ for $j=1,2$. The pricing policy consists of an exploration phase and an exploitation phase. During exploration, the seller uses a uniform pricing policy to collect data for estimating the demand parameters $\bm{\theta}$, resulting in estimates $\widehat{\bm{\theta}}$. In the exploitation phase, the seller applies the fairness-aware pricing policy: $p_{j}(\widehat{\bm{\theta}},\bm{x}_{t},\delta^{\prime})$ for $j=0,1$, where $\delta^{\prime}$ reflects the fairness constraint used in the algorithm. The pricing policy $p_{j}(\widehat{\bm{\theta}},\bm{x}_{t},\delta^{\prime})$ can be viewed as a plug-in estimator of $p^{*}_{j}(\bm{\theta},\bm{x}_{t},\delta)$, where $\widehat{\bm{\theta}}$ is obtained from the exploration phase. The fairness level $\delta^{\prime}$ used in the algorithm is typically chosen close to the global fairness parameter $\delta$.
We conduct new experiments to evaluate the performance of the proposed pricing policy under the one-time manipulation cost setting. We set $C_{0}=0.8,\delta=0.799$, and the length of the exploration phase as $T_{0}=\lceil 10\sqrt{T}\rceil$, with a total time horizon of $T=10000$. Following Algorithm 1, we set $\delta^{\prime}=\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}$, and vary the adjustment parameter $c_{\delta}=1,2,3$. Figure 12 presents the regret plots for our policy compared to the benchmark. In all cases, our proposed policy achieves substantially lower regret than the benchmark, demonstrating its effectiveness even when the cost of strategic manipulation is one-time fixed.
Refer to caption Refer to caption Refer to caption
Figure 12: Regret plots for the two policies. The three subplots show the regrets of three different scenarios, $c_{\delta}\in\{1,2,3\}$.
Appendix S.7 Discussion on Lower Bound
We derived a lower bound under the non-contextual pricing setting in Theorem 3. In this section, we present a lower bound in the contextual pricing setting under stronger assumptions by following Chai et al. (2024). We construct a special instance that satisfies Theorem 3 in Chai et al. (2024), which allows us to invoke their result establishing a dimension-free $\Omega(\sqrt{T})$ regret lower bound in contextual pricing. We include their assumption below for clarity:
Assumption 4 (Chai et al. (2024)).
  1. 1.
    The distribution $\mu$ of feature vectors $\bm{x}\in\mathbb{R}^{d}$ satisfies an anti-concentration property: there exists a constant $c_{\mathrm{ac}}>0$ such that for any symmetric matrix $\bm{A}\in\mathbb{S}^{d}$ with $\mathrm{rank}(\bm{A})\leq 4$, $\mathbb{E}_{\bm{x}\sim\mu}\left[(\bm{x}^{\top}A\bm{x})^{2}\right]\geq c_{\mathrm{ac}}\cdot\|\bm{A}\|_{F}^{2}.$
  2. 2.
    The demand model is non-degenerate, meaning there exists a constant $0<c<1$ such that $\|\bm{\alpha}_{j}\|_{2}\geq c,\quad\|\bm{\beta}_{j}\|_{2}\geq c,\quad\text{and}\quad\left|\bm{\alpha}_{j}^{\top}\bm{\beta}_{j}\right|\leq(1-c)\|\bm{\alpha}_{j}\|_{2}\cdot\|\bm{\beta}_{j}\|_{2}$ for $j=0,1$.
In our construction, we assume the following demand model:
$y_{jt}=\bm{\beta}_{j}^{\top}\widetilde{\bm{x}}_{t}+p_{t}\cdot(\widetilde{\bm{x}}_{t}^{\top}\bm{\alpha}_{j})+\epsilon_{t},$
where $\widetilde{\bm{x}}_{t}=(1,\bm{x}_{t}^{\top})^{\top}\in\mathbb{R}^{d+1}$ and $\epsilon_{t}$ is from the normal distribution $N(0,\sigma_{\epsilon}^{2})$. We set $\bm{\alpha}_{j}=(\alpha_{j},0,\dots,0)^{\top}\in\mathbb{R}^{d+1}$. Suppose that buyers can perfectly learn the price difference, and Assumption 4 holds. We further assume that the fairness constraint is not binding. Under this setup, the problem instance aligns with the setting of Theorem 3 in Chai et al. (2024), and therefore a dimension-free regret lower bound of $\Omega(\sqrt{T})$ applies.
Appendix S.8 Framework of a Nash Equilibrium Model
[…]